OpenStudy (anonymous):

I need help finding the first 5 terms of the sequence: an =n^2-2n

OpenStudy (jdoe0001):

\(\large{ \bf a_n =n^2-2n\\ \quad \\ a_{\color{red}{ 1}} ={\color{red}{ 1}}^2-2\cdot {\color{red}{ 1}}\qquad a_{\color{red}{ 2}} ={\color{red}{ 2}}^2-2\cdot {\color{red}{ 2}}\qquad a_{\color{red}{ 3}} ={\color{red}{ 3}}^2-2\cdot {\color{red}{ 3}}\\ \quad \\ a_{\color{red}{ 4}} ={\color{red}{ 4}}^2-2\cdot {\color{red}{ 4}}\qquad a_{\color{red}{ 5}} ={\color{red}{ 5}}^2-2\cdot {\color{red}{ 5}}}\)